OpenStudy (anonymous):

limit x-->0, [2tan^2(x)]/ x^2 find the limit

OpenStudy (sirm3d):

\[\large \lim_{x \rightarrow 0}\frac{2\tan^2x}{x^2}=2\lim_{x \rightarrow 0}\frac{\sin^2x}{x^2 \cos^2x}=2\lim_{x \rightarrow 0}\frac{\sin^2x}{x^2}\lim_{x \rightarrow 0}\frac{1}{\cos^2x}=2(1)(1)=2\]

OpenStudy (anonymous):

You're a lifesaver! Thanks sirm3d.